OpenStudy (anonymous):

Simplify this equation and state any restrictions on the variables. (x+1)/((x^2)+2x-8)-(x)/(4x-8)

OpenStudy (anonymous):

\[\frac{ x+1 }{ x ^{2}+2x-8 }- \frac{ x }{4x-8 }\]

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this?

jimthompson5910 (jim_thompson5910):

x^2 + 2x - 8 factors to ???

OpenStudy (anonymous):

i don't remember how to factor

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Maaf1234 @Dodo1 @zepdrix @mathstudent55

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to -8 and add to 2 at the same time these two numbers are -2 and 4

jimthompson5910 (jim_thompson5910):

so it factors to (x-2)(x+4)

jimthompson5910 (jim_thompson5910):

4x-8 factors to 4(x-2)

jimthompson5910 (jim_thompson5910):

this means \[\large \frac{ x+1 }{ x^2+2x-8 }- \frac{ x }{4x-8 }\] turns into \[\large \frac{ x+1 }{ (x-2)(x+4) }- \frac{ x }{ 4(x-2) }\]

jimthompson5910 (jim_thompson5910):

now you need to get the denominators the same

jimthompson5910 (jim_thompson5910):

the LCD is 4(x-2)(x+4) \[\large \frac{ x+1 }{ (x-2)(x+4) }- \frac{ x }{ 4(x-2) }\] \[\large \frac{ 4(x+1) }{ 4(x-2)(x+4) }- \frac{ x }{ 4(x-2) }\] \[\large \frac{ 4(x+1) }{ 4(x-2)(x+4) }- \frac{ x(x+4) }{ 4(x-2)(x+4) }\] \[\large \frac{ 4(x+1) - x(x+4) }{ 4(x-2)(x+4) }\] I'll let you finish

OpenStudy (anonymous):

@jim_thompson5910 so would it be \[\frac{ 6x+4 }{ 4x+4 }\]

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

\[\large \frac{ 4(x+1) - x(x+4) }{ 4(x-2)(x+4) }\] \[\large \frac{ 4x+4 - x^2-4x }{ 4(x-2)(x+4) }\] \[\large \frac{ 4 - x^2 }{ 4(x-2)(x+4) }\] keep going

OpenStudy (anonymous):

\[\frac{ 4-x^{2} }{ 4x-4}\]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\large \frac{ 4 - x^2 }{ 4(x-2)(x+4) }\] \[\large \frac{ -(x-2)(x+2) }{ 4(x-2)(x+4) }\] \[\large \frac{ -(x+2) }{ 4(x+4) }\] \[\large -\frac{ x+2 }{ 4x+16 }\]

OpenStudy (anonymous):

Omg so my answer is wrong smh

jimthompson5910 (jim_thompson5910):

well just keep practicing and you'll get better

OpenStudy (anonymous):

Is this it or it there more?

jimthompson5910 (jim_thompson5910):

nope that's it

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw